9. Suppose of we have 2 proposals (A and B) available for investment. For proposal A we need to invest S10,000 today and we will receive S12,000 in 1 year. For proposal B, we need to invest $11,000 today and we will receive $13,125 in 1 year. This information can be seen on the Table below Assume that we will only invest our money if we have a return at least greater than $100. Year A 0 -S10,000-$11,000 $12,000 $13,125 (a) Assuming we have a TVOM of 1090, which proposal should we choose and how much will be our return? Please, specify if you are using Rule#3 or Rule#4 of DCF. (b) Assuming now we have a TVOM of 1500, which proposal should we choose and how much will be our return? Please, specify if you are using Rule#3 or Rule#4 of DCF. (c) Assuming now we have a TVOM of 2090, which proposal should we choose and how much will be our return? Please, specify if you are using Rule#3 or Rule#4 of DCF.
Please use rule 3 not 4
Answer #1

Future value = Present value (1+r)n

For Investment A

12000 = 10,000(1+r)n

1+r = 1.2

r = 20%

Similarly,

13125 = 11000(1+r)

1+r = 1.19

r = 19%

From the above calculation we have seen that proposal A having 20% return than compare to proposal B and proposal A was receiving more return than in TVOM.

So in all cases it is suggested to invest in A.